The reaction 2A → A2 was experimentally determined to be second order with a rate constant, k, equal to 0.0265 M–1min–1. If the initial concentration of A was 3.00 M, what was the concentration of A (in M) after 180.0 min?

  M ?
Answer #1

Given, the balanced second order reaction,

2A A2

Also given,

Rate constant(k) = 0.0265 M-1 min-1

The initial concentration of A (No) = 3.00 M

Time (t) = 180.0 min

Nt = ? The concentration of A(in M) after time "t".

We know, The second order rate law equation,

1 / [Nt] = kt + 1 / [No]

Substituting the known values in the formula,

1 / [Nt] = 0.0265 M-1 min-1 x 180.0 min. + 1 / 3.00 M

1 / [Nt] = 4.77 + 0.3333

1 / [Nt] = 5.1033

[Nt] = 1 / 5.1033

[Nt] = 0.196 M